# A Survey on Diffusion Models for Recommender Systems

### v1

@@ Abstract.
The rapid advancement of online services has positioned recommender systems (RSs) as crucial tools for mitigating information overload and delivering personalized content across e-commerce, entertainment, and social media platforms. While traditional recommendation techniques have made significant strides in the past decades, they still suffer from limited generalization performance caused by factors like inadequate collaborative signals, weak latent representations, and noisy data. In response, diffusion models (DMs) have emerged as promising solutions for recommender systems due to their robust generative capabilities, solid theoretical foundations, and improved training stability compared with other generative modeling techniques like variational autoencoders (VAEs) or generative adversarial networks (GANs). To this end, in this paper, we present the first comprehensive survey on diffusion models for recommendation, and draw a bird’s-eye view from the perspective of the whole pipeline in real-world recommender systems. We systematically categorize existing research works into three primary domains: (1) diffusion for data engineering & encoding, focusing on data augmentation and representation enhancement; (2) diffusion as recommender models, employing diffusion models to directly estimate user preferences and rank items; and (3) diffusion for content presentation, utilizing diffusion models to generate personalized content such as fashion and advertisement creatives. Our taxonomy highlights the unique strengths of diffusion models in capturing complex data distributions and generating high-quality, diverse samples that closely align with user preferences. We also summarize the core characteristics of the adapting diffusion models for recommendation, and further identify key areas for future exploration, which helps establish a roadmap for researchers and practitioners seeking to advance recommender systems through the innovative application of diffusion models. To further facilitate the research community of recommender systems based on diffusion models, we actively maintain a GitHub repository for papers and other related resources in this rising direction [^1] https://github.com/CHIANGEL/Awesome-Diffusion-for-RecSys. footnotetext: $\dagger$ Weinan Zhang is the corresponding author.
@@ 1. Introduction
As illustrated in Figure 1, in the past decades, we have witnessed significant progress in the research on recommender systems, shifting from traditional techniques like collaborative filtering (CF) (He et al., 2016) to more advanced deep learning methodologies (Lin et al., 2023b). However, they usually suffer from limited generalization performance on account of inadequate collaborative signals (Lin et al., 2023a), weak latent representations (Du et al., 2024), noisy data scenarios (Wang et al., 2021a). Therefore, the generative models, e.g., variational autoencoders (VAEs) (Kingma and Welling, 2014; Rezende et al., 2014) and generative adversarial networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Brock et al., 2019), turn out a promising solution to mitigate the above challenges for recommendation due to their generative nature and solid theoretical foundation. However, these models still have their own limitations such as restricted representation capability (Wang et al., 2023b) and training instability (Becker et al., 2022).
Figure 1. (a) The development trends and representational works of recommendation methods from traditional collaborative filtering (CF) based methods to generative methods, i.e., autoencoder (AE), variational autoencoder (VAE), generative adversarial network (GAN), and diffusion model. (b) The cumulative paper count of diffusion-based recommendation methods according to the timeline. (c) The paper distribution of diffusion-based recommendation methods over venues.
Recently, diffusion models (DMs) (Ho et al., 2020; Song et al., 2020) have emerged as the state-of-the-art of generative modeling, achieving substantial success in various domains including computer vision (Lugmayr et al., 2022), audio generation (Lee and Han, 2021), natural language processing (Austin et al., 2021), and reinforcement learning (Zhu and Zhao, 2023). Unlike other earlier generative models like VAEs and GANs, diffusion models leverage a denoising framework to effectively reverse a multi-step noising process to generate synthetic data that aligns closely with the distribution of the training data. This ensures the remarkable capabilities of diffusion models in capturing the multi-grained feature representations and generating high-quality samples, as well as maintaining enhanced training stability. As a consequence, as shown in Figure 1, an increasing range of pioneer attempts has been made to employ diffusion models for recommendation, achieving notable progress in boosting the performance of different canonical recommendation processes, e.g., data augmentation (Wang et al., 2024b), user modeling (Zhao et al., 2024a), and content personalization (Yang et al., 2024d), etc. We summarize the three main characteristics that make diffusion models attractive in the context of recommender systems:
@@
    Superior representation learning. Diffusion models are known for their remarkable abilities to learn high-quality, low-dimensional representations of the source data in a probabilistic generative manner (Yang et al., 2024b; Fuest et al., 2024). In the context of recommender systems, they can effectively capture the underlying latent factors and representations based on multi-modal data (e.g., interaction data, complex user behaviors, and item attributes), thus making more accurate predictions about user preferences and generating recommendations tailored to individual tastes.
@@
    Flexible internal structure. As a general learning framework for generative modeling, the internal structures of diffusion models (i.e., the backbone design) are fairly flexible, and can be integrated with other deep learning models like U-Net (Ho et al., 2020) and Transformers (Peebles and Xie, 2023). This allows for flexible backbone designs of diffusion models to effectively incorporate different types of heterogeneous information (e.g.user demographics, temporal dynamics, and contextual cues) into the recommendation process.
@@
Given the advantages of applying diffusion models to recommendation, as well as the proliferation of research in the community, we believe that the time is right to conduct a comprehensive survey to systematically summarize the current research progress and provide inspirations for the future exploration, in terms of the adaption of diffusion models in recommender systems.
@@
Diffusion models are closely related to generative modeling and self-supervised learning. There exist several related survey works that delve into the potential of generative models (Li et al., 2023b; Deldjoo et al., 2024; Li et al., 2024d; Xu et al., 2024b; Liang et al., 2024; Deldjoo et al., 2021; Zhang et al., 2020b) or self-supervised learning techniques (Jing et al., 2023; Liu et al., 2023c; Yu et al., 2023b) for recommender systems. For example, Liu et al. (2023c) and Yu et al. (2023b) conduct reviews on the self-supervised learning (SSL) based techniques for recommendation like contrastive learning and sequence modeling. Deldjoo et al. (2021) centers on adversarial recommender systems, where the generative adversarial networks (GANs) are widely employed for security and robustness. Liang et al. (2024) investigate the applications of variational autoencoders (VAEs) in recommender systems based on their generative Bayesian nature. There is also a range of surveys (Li et al., 2023b; Deldjoo et al., 2024; Li et al., 2024d; Xu et al., 2024b) focus on the generative recommendation with the assistance of large language models (LLMs), which serve as the most popular foundation models for various downstream scenarios in the past years. However, none of these surveys concentrate on the applications of diffusion models for recommendation. There still lacks a bird’s-eye view of how recommender systems can embrace diffusion models and integrate them into different parts of the recommendation pipeline, which is essential in building a technical roadmap to systematically guide the research, as well as industrial practice, of recommender systems empowered by diffusion models.
@@
To this end, in this paper, we aim to conduct a timely and comprehensive survey on the adaption of diffusion models to recommender systems. As depicted in Figure 3, we analyze the latest research progress and categorize the existing works according to different roles that diffusion models play at different parts of the modern deep learning based recommender system pipeline:
@@
  • Diffusion for data engineering & encoding. Data engineering & encoding is generally referred to as the process of manipulating and transforming the raw data collected online into structured data or neural embeddings for the downstream recommenders. As a powerful class of generative models, diffusion models have shown remarkable capabilities in both data augmentation and representation enhancement, both of which help improve the downstream recommendation performance.
@@
  • Diffusion as recommender model. The recommender aims to estimate a given user’s preference towards each candidate item, and finally arrange a ranked list of items presented to the user. According to different types of tasks the recommender aims to solve, we classify the diffusion-model-based (DM-based) recommenders into three categories: collaborative recommendation, context-aware recommendation, and cross-domain recommendation.
@@
  • Diffusion for content presentation. Equipped with diffusion models, we can move one step further from personalized recommendation to individualized content generation. That is, every single item can obtain different presentation contents (e.g., creatives, thumbnails) produced by diffusion models for different users or groups, which can largely promote the user satisfaction.
@@
  • Unified and structured taxonomy. We introduce a well-organized categorization to classify the existing research works into three major types according to the different roles the diffusion models play: diffusion for data engineering & encoding, diffusion as recommender model, and diffusion for content presentation. This taxonomy provides the readers with a coherent roadmap, and helps recognize the trend in applications of diffusion models to recommender systems from multiple perspectives.
@@ 2. Preliminary
Figure 2. (a) The illustration of a deep learning based recommender system pipeline, which is characterized by three major stages: data engineering & encoding, recommender model, and content presentation. (b) An overview of diffusion models for data analysis and generation with various modalities.
@@ 2.1. Modern Recommender Systems
As shown in Figure 2(a), there are generally three key components in deep learning based recommender systems:
@@
  • Data engineering & encoding. As the foundational pillar of modern deep learning based recommender systems, data engineering & encoding consists of two primary processes: (1) data engineering and (2) data encoding. Data engineering generally encompasses selecting, manipulating, transforming, and augmenting the raw data collected online into structured data that is suitable as inputs of neural recommendation models. It does not only involve feature-level manipulation, but also includes sample-level augmentation and synthesis. The outputs of data engineering possess various forms of features with different modalities, e.g., IDs, texts, images, audio, etc. The data encoding process takes as input the processed structured data, and produces the corresponding neural embeddings for the downstream deep recommender models. Various encoders are employed depending on the data modality. Typically, this process is designed as an embedding layer for one-hot encoded categorical features for traditional recommender models. Features of other modalities further require different encoders for the data encoding process, e.g., vision models for visual features, language models for textual features.
@@
  • Recommender Model. The deep recommender model is the core algorithmic engine of a recommender system, tasked with selecting or ranking the top-relevant items to satisfy users’ information needs based on the structured data or neural embeddings produced by the data engineering & encoding stage. Researchers develop a variety of neural methods to accurately estimate the user interests and behavior patterns based on various techniques like sequence modeling (Cheng et al., 2024; Liu et al., 2023a, 2024a) and graph neural networks (Wang et al., 2018, 2019). According to the task types and input data formats, the design of deep recommender models can be generally classified into three categories: (1) collaborative recommendation, (2) context-aware recommendation, and (3) cross-domain recommendation [^2] We categorize multi-domain or multi-scenario recommendation as cross-domain recommendation as well.. The collaborative recommendation is generally referred to as the collaborative filtering (CF) based methods simply based on the user-item co-occurrence matrix. The context-aware recommendation takes into account the contextual information surrounding a user’s request to provide more precise suggestions, e.g., user behavior sequence for user profiling, or knowledge graph for item content understanding. The cross-domain recommendation further introduces transfer learning among multiple domains or scenarios to enhance the recommendation performance, which is particularly helpful to overcome the data sparsity issue.
@@
  • Content presentation. After arranging the ranked list of items with the recommender model, the content presentation serves as the final touch that brings the recommendations to life, ensuring the recommended items are delivered in a manner that is engaging, accessible, and contextually relevant to the target user. Conventionally, it involves the strategic placement of recommended items, the manual usage of visual elements such as images, videos, and text to attract user attention, and the implementation of interactive features that encourage users to be more willing to explore and interact with the displayed items. Nowadays, with the rapid development of generative models (e.g., large language models (Zhao et al., 2023), diffusion models (Ho et al., 2020)), we can take one step further from personalized recommendation to individualized content generation. That is, the concrete content of each recommended item (e.g., creatives, titles, thumbnails) can be dynamically edited or generated based on user profiles, device types, and environmental contexts, thereby increasing the likelihood of user satisfaction and conversion.
@@ 2.2. Foundations of Diffusion Models
Typically, as illustrated in Figure 2(b), the training procedure of diffusion models includes two stages: the forward process (diffusion) and the reverse process (denoising). In the forward diffusion process, the model turns a data sample into pure random noise by incrementally adding noise for multiple steps, which is usually a Markov process with each step depending only on the preceding one. Then, the reverse denoising process learns to remove the noise to reconstruct the original data sample, essentially reversing the forward process. In this way, the model learns to remove the noise added during the diffusion process, and thereby generates samples from the same distribution as the training data. According to different designs of the forward and backward processes, the common frameworks for diffusion models include denoising diffusion probabilistic models (Ho et al., 2020), score-based generative methods (Song et al., 2020), conditional diffusion models (Rombach et al., 2022b), etc. Next, we will provide a concise introduction to these three classical types of diffusion models.
@@ 2.2.1. Denoising Diffusion Probabilistic Models (DDPMs)
Forward Diffusion Process. Assume that there is an initial clean data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ sampled from a given data distribution $q(\mathbf{x})$. The ensuing forward diffusion process adulterates the initial data $\mathbf{x}_{0}$ by incrementally superimposing Gaussian noise, ultimately aiming to progress towards convergence with the standard Gaussian distribution (i.e., pure noise). During the forward process with maximum steps up to $K$, we will materialize a sequence of distributed latent data $[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{K}]$, which can be formulated as a Markov chain transforming from $\mathbf{x}_{k-1}$ to $\mathbf{x}_{k}$ with a diffusion transition kernel:
@@
where $\beta_{k}\in(0,1)$ serves as a variance schedule to control the step size, $\mathbf{I}$ is the identity matrix with the same dimension as the input data $\mathbf{x}_{k-1}$, and $\mathcal{N}(\mathbf{x};\mu,\sigma)$ is a Gaussian distribution of $\mathbf{x}$ with the mean $\mu$ and the standard deviation $\sigma$. According to the property of the Gaussian kernel, we can get $\mathbf{x}_{k}$ directly from $\mathbf{x}_{0}$ by applying a series of transition kernels of Eq. 2:
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is the Gaussian noise. Specifically, it is designed $\bar{\alpha}_{K}\approx 0$ so that
@@
That is, the reverse denoising process can start with any Gaussian noise. To sum up, the forward diffusion process gradually injects noise into the initial data $\mathbf{x}_{0}$ until it nearly aligns with the standard Gaussian distribution.
@@
where the mean $\mu_{\theta}(\cdot)$ and variance $\sigma_{\theta}(\cdot)$ are parameterized by $\theta$. The model aims to learn the data distribution via the model distribution ${p}_{\theta}(\mathbf{x}_{0})$ during the reverse denoising process.
Training. In order to approximate the ground-truth data distribution $q(\mathbf{x}_{0})$, DDPM is trained to minimize the variational upper bound on the negative log-likelihood (NLL):
@@
where $\alpha_{k}=1-\beta_{k}$ and $\bar{\alpha}_{k}=\prod_{k=1}^{K}\alpha_{k}$. In this way, $L_{k-1}$ can be equally regarded as the expected value of the L2 loss between the two mean coefficients:
@@
(11) $\mathbb{E}_{k\sim\mathcal{U}(1,K),\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg{[}{\lambda(k)\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{k},k)\right\|^{2}_{2}}\bigg{]},$
@@
(12) $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})$ $=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\sigma_{\theta}(\mathbf{x}_{k},k)\mathbf{I})$
@@
where $z\sim\mathcal{N}(0,\mathbf{I})$, and $\beta_{k}\approx\sigma_{\theta}^{2}(\mathbf{x}_{k},k)$. While the vanilla DDPM (Ho et al., 2020) can only be applied to continuous data like audio (Lee and Han, 2021) and image (Lugmayr et al., 2022), researchers have extended its applications to various discrete data like categorical data (Hoogeboom et al., 2021), tabular data (Kotelnikov et al., 2023), and textual data (Zhu and Zhao, 2023), etc.

### v2

@@ Abstract.
The rapid advancement of online services has positioned recommender systems (RSs) as crucial tools for mitigating information overload and delivering personalized content across e-commerce, entertainment, and social media platforms. While traditional recommendation techniques have made significant strides in the past decades, they still suffer from limited generalization performance caused by factors like inadequate collaborative signals, weak latent representations, and noisy data. In response, diffusion models (DMs) have emerged as promising solutions for recommender systems due to their distinguished generative capability, superior representation learning, and flexible internal structure. To this end, in this paper, we present the first comprehensive survey on diffusion models for recommendation, and draw a bird’s-eye view from the perspective of the whole pipeline in real-world recommender systems. We systematically categorize existing research works into three primary domains: (1) diffusion for data engineering & encoding that focuses on data augmentation and representation enhancement; (2) diffusion as recommendation models to directly estimate user preferences and rank items; and (3) diffusion for content presentation to generate personalized content such as fashion and advertisement creatives. Our taxonomy highlights the unique strengths of diffusion models in capturing complex data distributions and generating high-quality, diverse samples that closely align with user preferences. We also summarize the core characteristics of adapting diffusion models for recommendation, and further identify key areas for future exploration, which helps establish a roadmap for researchers and practitioners seeking to advance recommender systems through the innovative application of diffusion models. To further facilitate the research community of recommender systems based on diffusion models, we actively maintain a GitHub repository for papers and other related resources in this rising direction [^1] https://github.com/CHIANGEL/Awesome-Diffusion-for-RecSys. footnotetext: $\dagger$ Weinan Zhang is the corresponding author.
@@ 1. Introduction
As illustrated in Figure 1, in the past decades, we have witnessed significant progress in the research on recommender systems, shifting from traditional techniques like collaborative filtering (CF) (He et al., 2016) to more advanced deep learning methodologies (Lin et al., 2023b). However, they usually suffer from limited generalization performance on account of inadequate collaborative signals (Lin et al., 2023a), weak latent representations (Du et al., 2024), noisy data scenarios (Wang et al., 2021a). Therefore, the generative models, e.g., variational autoencoders (VAEs) (Kingma and Welling, 2014; Rezende et al., 2014) and generative adversarial networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Brock et al., 2019), turn out a promising solution to mitigate the above challenges for recommendation due to their generative nature and solid theoretical foundation. The self-supervised characteristic enables them to learn the non-linear probabilistic latent space and thereby precisely capture the users’ dynamic and ever-evolving preferences (Fan et al., 2022b; Liang et al., 2024). However, these models still have their own limitations such as restricted representation capability (Wang et al., 2023b) and training instability (Becker et al., 2022).
Figure 1. (a) The development trends and representational works of recommendation methods from traditional collaborative filtering (CF) based methods to generative methods, i.e., autoencoder (AE), variational autoencoder (VAE), generative adversarial network (GAN), and diffusion model. (b) The cumulative paper count of diffusion-based recommendation methods according to the timeline. (c) The paper distribution of diffusion-based recommendation methods over venues, where the venue name is followed by the exact number of published papers.
Recently, diffusion models (DMs) (Ho et al., 2020; Song et al., 2020) have emerged as the state-of-the-art of generative modeling paradigms, achieving substantial success in various domains including computer vision (Lugmayr et al., 2022), audio generation (Lee and Han, 2021), natural language processing (Austin et al., 2021), and reinforcement learning (Zhu and Zhao, 2023). Unlike other earlier generative models such as VAEs and GANs, diffusion models leverage a denoising framework that effectively reverses a multi-step noising process to generate synthetic data that aligns closely with the distribution of the training data. This ensures the remarkable capabilities of diffusion models in capturing the multi-grained feature representations and generating high-quality samples, as well as maintaining enhanced training stability. We summarize the three main characteristics that make diffusion models attractive in the context of recommender systems:
@@
    Superior representation learning. Diffusion models are powerful self-supervised methods, and are known for their remarkable abilities to learn high-quality, low-dimensional representations of the source data in a probabilistic generative manner (Yang et al., 2024b; Fuest et al., 2024). In the context of recommender systems, they can effectively capture the underlying latent factors and representations based on multi-modal data (e.g., interaction data, complex user behaviors, and item attributes), thus making more accurate predictions about user preferences and generating recommendations tailored to individual tastes.
@@
    Flexible backbone structure. As a general learning framework for generative modeling, the internal structures of diffusion models (i.e., the backbone design) are fairly flexible, and can be integrated with other deep learning models like U-Net (Ho et al., 2020) and Transformers (Peebles and Xie, 2023). This allows for flexible backbone designs of diffusion models to effectively incorporate different types of heterogeneous information (e.g.user demographics, temporal dynamics, and contextual cues) into the recommendation process.
@@
As a consequence, as shown in Figure 1, an increasing range of pioneer attempts has been made to employ diffusion models for recommendation, achieving notable progress in boosting the performance of different canonical recommendation processes, e.g., data augmentation (Wang et al., 2024b), user modeling (Zhao et al., 2024a), and content personalization (Yang et al., 2024d), etc. Given the advantages of applying diffusion models to recommendation, as well as the proliferation of research in the community, we believe that the time is right to conduct a comprehensive survey to systematically summarize the current research progress and provide inspirations for the future exploration, in terms of the adaption of diffusion models in recommender systems.
@@
Diffusion models are closely related to generative modeling and self-supervised learning. There exist several related survey works that delve into the potential of generative models (Li et al., 2023b; Deldjoo et al., 2024; Li et al., 2024d; Xu et al., 2024b; Liang et al., 2024; Deldjoo et al., 2021; Zhang et al., 2020b) or self-supervised learning techniques (Jing et al., 2023; Liu et al., 2023c; Yu et al., 2023b) for recommender systems. For example, Liu et al. (2023c) and Yu et al. (2023b) conduct reviews on the self-supervised learning (SSL) based techniques for recommendation like contrastive learning and sequence modeling. Deldjoo et al. (2021) centers on adversarial recommender systems, where generative adversarial networks (GANs) are widely employed for security and robustness. Liang et al. (2024) investigate the applications of variational autoencoders (VAEs) in recommender systems based on their generative Bayesian nature. There is also a range of surveys (Li et al., 2023b; Deldjoo et al., 2024; Li et al., 2024d; Xu et al., 2024b) focus on the generative recommendation with the assistance of large language models (LLMs), which serve as the most popular foundation models for various downstream scenarios in the past years. However, none of these surveys concentrate on the applications of diffusion models for recommendation. There still lacks a bird’s-eye view of how recommender systems can embrace diffusion models and integrate them into different parts of the recommendation pipeline, which is essential in building a technical roadmap to systematically guide the research and industrial practice, of recommender systems empowered by diffusion models.
@@
To this end, in this paper, we aim to conduct a timely and comprehensive survey on the adaption of diffusion models to recommender systems. As depicted in Figure 2(a), we analyze the latest research progress and categorize the existing works according to different roles that diffusion models play in the modern recommender system pipeline:
@@
  • Diffusion for data engineering & encoding. Data engineering & encoding is generally referred to as the process of manipulating and transforming the raw data collected online into structured data or neural embeddings for the downstream recommenders. As a powerful class of generative models as well as self-supervised learning algorithms, diffusion models have shown remarkable capabilities in both data augmentation and representation enhancement, both of which help improve the downstream recommendation performance.
@@
  • Diffusion as recommendation model. The recommendation model aims to estimate a given user’s preference towards each candidate item, and finally arrange a ranked list of items presented to the user. According to different types of tasks the recommender aims to solve, we classify the diffusion-based recommenders into three categories: collaborative recommendation, context-aware recommendation, and other applications.
@@
  • Diffusion for content presentation. Equipped with powerful generative capabilities of diffusion models, we can move one step further from personalized recommendation to individualized content generation. That is, every single item can obtain different presentation contents (e.g., creatives, thumbnails) produced by diffusion models for different users or groups, which can largely promote the user satisfaction. We categorize the research works in this line based on the different types of content to be generated: fashion generation, ad creative generation, and general content generation.
@@
  • Unified and structured taxonomy. We introduce a well-organized categorization to classify the existing research works into three major types according to the different roles the diffusion models play: diffusion for data engineering & encoding, diffusion as recommendation model, and diffusion for content presentation. This taxonomy provides the readers with a coherent roadmap, and helps recognize the trend in applications of diffusion models to recommender systems from multiple perspectives.
@@ 2. Preliminary
Figure 2. (a) The illustration of a deep learning based recommender system pipeline, which is characterized by three major stages: data engineering & encoding, recommendation model, and content presentation. (b) An overview of diffusion models for data analysis and generation via the diffusion-denoising process.
@@ 2.1. Modern Recommender Systems
As shown in Figure 2(a), there are generally three key components in deep learning based recommender system pipelines. We briefly introduce each of them as follows:
@@
  • Data engineering & encoding. The data pre-processing stage serves as the foundational pillar of modern deep learning based recommender systems, and generally consists of two primary processes: (1) data engineering and (2) data encoding. Data engineering generally encompasses selecting, manipulating, transforming, and augmenting the raw data collected online into structured data that is suitable as inputs of neural recommendation models, which can reveal the important signals and patterns for user modeling. It does not only involve feature-level editing (e.g., cross-feature generation), but also includes sample-level synthesis (e.g., user interaction records generation and negative sample mining). The structured outputs of data engineering possess various forms of features with different modalities, e.g., IDs, texts, images, audio, etc. The data encoding process takes as input the processed structured data from data engineering, and produces the corresponding neural embeddings for the downstream deep recommendation models. Various encoders are employed depending on the data modality, e.g., vision models for visual features, language models for textual features.
@@
  • Recommendation Model. The deep recommendation model is the core algorithmic engine of a recommender system, tasked with selecting or ranking the top-relevant items to satisfy users’ information needs based on the structured data or neural embeddings produced by the data engineering & encoding stage. Researchers develop a variety of neural methods to accurately estimate the user interests and behavior patterns based on various techniques like sequence modeling (Cheng et al., 2024; Liu et al., 2023a, 2024a) and graph neural networks (Wang et al., 2018, 2019). According to the task types and input data formats, the design of deep recommendation models can be generally classified into three categories: (1) collaborative recommendation, (2) context-aware recommendation, and (3) other applications. The collaborative recommendation is generally referred to as the collaborative filtering (CF) based methods simply based on the user-item co-occurrence matrix (He et al., 2017). The context-aware recommendation takes into account the contextual information surrounding a user’s request to provide more precise recommendations, e.g., user behavior sequence for user profiling (Zhou et al., 2018), or knowledge graph for item content understanding (Wang et al., 2018). Finally, other applications refers to a broader range of tasks that are closely related to the aforementioned collaborative or context-aware recommendation, e.g., learning to rank (Cao et al., 2007), and bidding for online advertising (Cai et al., 2017).
@@
  • Content presentation. After arranging the ranked list of items with the recommendation model, the content presentation serves as the final touch that brings the recommendations to life, ensuring the recommended items are delivered in a manner that is engaging, accessible, and contextually relevant to the target user. Conventionally, it involves the strategic placement of recommended items, the manual usage of visual elements such as images, videos, and text to attract user attention, and the implementation of interactive features that encourage users to be more willing to explore and interact with the displayed items. Despite some appealing manual designs, the content of one certain item presented to different users remains uniform. Nowadays, with the rapid development of generative models (e.g., large language models (Zhao et al., 2023), diffusion models (Ho et al., 2020)), we can take one step further from unified content presentation to individualized content generation. That is, the concrete content of each recommended item (e.g., creatives, titles, thumbnails) can be dynamically edited or generated based on user profiles, device types, and environmental contexts, thereby increasing the likelihood of user attention and conversion.
@@ 2.2. Foundations of Diffusion Models
Typically, as illustrated in Figure 2(b), the training procedure of diffusion models includes two stages: the forward process (diffusion) and the reverse process (denoising). In the forward diffusion process, the model turns a data sample into pure random noise by incrementally adding noise for multiple steps, which is usually a Markov process with each step depending only on the preceding one. Then, the reverse denoising process learns to remove the noise to reconstruct the original data sample, essentially reversing the forward process. In this way, the model learns to remove the noise added during the diffusion process, and thereby generates samples from the same distribution as the training data. According to different designs of the forward and backward processes, the common frameworks for diffusion models include denoising diffusion probabilistic models (Ho et al., 2020), score-based generative methods (Song and Ermon, 2019; Song et al., 2020), conditional diffusion models (Rombach et al., 2022b), etc. Next, we will provide a concise introduction to these three classical types of diffusion models.
@@ 2.2.1. Denoising Diffusion Probabilistic Models (DDPMs)
Forward Diffusion Process. Assume that there is an initial clean data $\mathbf{x}_{0}\sim q(\mathbf{x})$ sampled from a given data distribution $q(\mathbf{x})$. The ensuing forward diffusion process adulterates the initial data $\mathbf{x}_{0}$ by incrementally superimposing Gaussian noise, ultimately aiming to progress towards convergence with the standard Gaussian distribution (i.e., pure noise). During the forward process with maximum steps up to $K$, we will materialize a sequence of distributed latent data $[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{K}]$, which can be formulated as a Markov chain transforming from $\mathbf{x}_{k-1}$ to $\mathbf{x}_{k}$ with a diffusion transition kernel:
@@
where $\beta_{k}\in(0,1)$ serves as a variance schedule to control the step size, $\mathbf{I}$ is the identity matrix with the same dimension as the input data $\mathbf{x}_{k-1}$, and $\mathcal{N}(\mathbf{x};\boldsymbol{\mu},\sigma\mathbf{I})$ is a Gaussian distribution of $\mathbf{x}$ with the mean $\mathbf{\mu}$ and the standard deviation $\sigma\mathbf{I}$. According to the property of the Gaussian kernel, we can get $\mathbf{x}_{k}$ directly from $\mathbf{x}_{0}$ by applying a series of transition kernels of Eq. 2:
@@
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the Gaussian noise. Specifically, it is designed $\bar{\alpha}_{K}\approx 0$ so that
@@
To sum up, the forward diffusion process gradually injects noise into the initial data $\mathbf{x}_{0}$ until it nearly aligns with the standard Gaussian distribution, and the reverse denoising process, as shown in Eq. 5 can start with any Gaussian noise.
@@
where the mean $\boldsymbol{\mu}_{\theta}(\cdot)$ and variance $\sigma_{\theta}(\cdot)$ are parameterized by $\theta$. The goal is to learn and approximate the data distribution via the model distribution ${p}_{\theta}(\mathbf{x}_{0})$ during the reverse denoising process.
Training. In order to approximate the ground-truth data distribution $q(\mathbf{x})$, DDPM is trained to minimize the variational upper bound on the negative log-likelihood (NLL):
@@
where $\alpha_{k}=1-\beta_{k}$ and $\bar{\alpha}_{k}=\prod_{i=1}^{k}\alpha_{i}$. In this way, $L_{k-1}$ can be equally regarded as the expected value of the L2 loss between the two mean coefficients:
@@
(11) $\mathbb{E}_{k\sim\mathcal{U}(1,K),\mathbf{x}_{0}\sim q(\mathbf{x}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg{[}{\lambda(k)\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{k},k)\right\|^{2}_{2}}\bigg{]},$
@@
(12) $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})$ $=\mathcal{N}(\mathbf{x}_{k-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{k},k),\sigma_{\theta}(\mathbf{x}_{k},k)\mathbf{I})$
@@
where $z\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\beta_{k}\approx\sigma_{\theta}^{2}(\mathbf{x}_{k},k)$. While the vanilla DDPM (Ho et al., 2020) are proposed to handle continuous data like audio (Lee and Han, 2021) and image (Lugmayr et al., 2022), researchers have extended its applications to other data mining scenarios like tabular data (Kotelnikov et al., 2023) and textual data (Zhu and Zhao, 2023) with specially designed adaptations.